OpenStudy (anonymous):

factor -x^2-2x+15=0 PLZ HELP lol

OpenStudy (anonymous):

(x+5)(x-3)

OpenStudy (anonymous):

gracias!!!

OpenStudy (anonymous):

I know how to do it, im just lazy lolz

OpenStudy (anonymous):

looooooooooooool

OpenStudy (anonymous):

yeepers

OpenStudy (mertsj):

The only problem is that (x+5)(x-3)=x^2+2x-15 not -x^2-2x+15

OpenStudy (mertsj):

So perhaps you should adjust your answer.

OpenStudy (anonymous):

(x+3)(x-5) to get the signs right

OpenStudy (anonymous):

i've divide both side by -1

OpenStudy (mertsj):

(x+3)(x-5)=x^2-2x-15 not -x^2 -2x-15

OpenStudy (anonymous):

it's (x+5)(x-3) not (x+3)(x-5)

OpenStudy (anonymous):

oh ok gotcha

OpenStudy (anonymous):

wait... the x in the equation is negative, so wouldn't one of the x's be negative in the factored part?

OpenStudy (mertsj):

\[-x^2-2x+15=-1(x^2+2x-15)=-1(x+5)(x-3)\]

OpenStudy (mertsj):

So: -1(x+5)(x-3)=0